How many milliters of a 0.200 M HNO3 solution are needed to dissolve 0.750 g of Cu?
Answer #1

Reaction is

Cu + 4 HNO3 .............> Cu(NO3)2 + 2 NO2 + 2 H2O

One mole Cu react with 4 mole HNO3.

0.750 g of Cu = mass / molar mass = 0.750 g / 63.5 g / mole = 0.01118 mole.

thus

mole of HNO3 = 4 * 0.01118 mole.= 0.0472 mole.

we have

0.200 M HNO3.

volume of HNO3 = 1000 ml * 0.0472 / 0.200 = 236.2 ml (answer)